OpenStudy (anonymous):

Write an equation in point-slope form of the line that passes through the points (-3,-4) and (3,4). @Here_to_Help15

OpenStudy (jhannybean):

First you need to find your slope, m. \[\ m = \frac{rise}{run} = \frac{y_2 - y_1}{x_2- x_1} = \frac{4-(-4)}{3-(-3)} = \frac{8}{9}\]Then you use point-slope form to find your equation, then change it to slope-intercept form to find the equation of the line . Point-slope : \[\ y-y_1 = m(x-x_1) \]\[\ y- (-3)=\left(\frac{8}{9}\right)(x-(-4))\] Reduce that.

OpenStudy (jhannybean):

once it's reduced, write it in the form \(\ y = mx+b\)

OpenStudy (anonymous):

let me try

OpenStudy (anonymous):

do i multipy 4 on each side

OpenStudy (jhannybean):

Why are you multiplying 4 and where did you get that from?

OpenStudy (jhannybean):

First reduce the left hand side, what do you get? Left hand only.

OpenStudy (jhannybean):

as in , what is y- (-3) ?

OpenStudy (jhannybean):

is it y+3 or y-3?

OpenStudy (anonymous):

y+3

OpenStudy (jhannybean):

alright, and x-(-4) . Is this x+4 or x-4?

OpenStudy (anonymous):

x+4 @Jhannybean

OpenStudy (jhannybean):

Alright, so you have \[\ y+3 = \left(\frac{8}{9}\right)(x+4)\] distribute the 8/9.\[\ y+3 = \frac{8}{9}x + \frac{8}{9}\cdot 4\]\[\ y+3 = \frac{8}{9}x + \frac{8\cdot 4}{9}\] What is 8 x 4?

OpenStudy (anonymous):

32

OpenStudy (anonymous):

@Jhannybean

OpenStudy (jhannybean):

good, so you have \[\ y + 3 = \frac{8}{9}x + \frac{32}{9}\]

OpenStudy (jhannybean):

Now you're going to subtract 3 from both sides of the equation.\[\ y= \frac{8}{9}x + \frac{32}{9}-3\]Make 3 into a fraction with a common denominator of 9.

OpenStudy (jhannybean):

\[\ y= \frac{8}{9}x + \frac{32}{9} -\frac{27}{9}\] 32 - 27 = 5\[ y= \frac{8}{9}x + \frac{5}{9}\]

OpenStudy (anonymous):

ok so that is how you got that anwser

OpenStudy (anonymous):

@Jhannybean

OpenStudy (jhannybean):

Yes it is :)